# GST-UNet: Spatiotemporal Causal Inference with Time-Varying Confounders

### v1

@@ 5 Experiments
Additional details–including exact simulation parameters, model architecture and execution setups, hyperparameter selection strategies, and validation procedures–can be found in Appendix C. Replication code is available at https://github.com/moprescu/GSTUNet.
@@ 5.1 Synthetic Data
Table 1: Comparison of models across horizons ($\tau$). Values represent RMSE ± standard deviation across mean test trajectories. GST-UNet (ours) is highlighted in the last row. Bold values indicate the smallest number in each column for each horizon. Relative changes for GST-UNet compared to the best baseline are shown in green (improvement) or red (decrease).
$\tau$ Model $\beta_{1}=0.0$ $\beta_{1}=0.5$ $\beta_{1}=1.0$ $\beta_{1}=1.5$ $\beta_{1}=2.0$ $\beta_{1}=2.5$ $\beta_{1}=3.0$
5 UNet+ 0.51 ± 0.00 0.62 ± 0.01 0.84 ± 0.01 1.03 ± 0.01 1.10 ± 0.01 1.16 ± 0.01 1.25 ± 0.01
STCINet 0.52 ± 0.00 0.68 ± 0.01 0.93 ± 0.01 1.11 ± 0.01 1.20 ± 0.01 1.33 ± 0.01 1.32 ± 0.01
GST-UNet 0.55 ± 0.01 0.61 ± 0.01 0.60 ± 0.01 0.61 ± 0.01 0.64 ± 0.01 0.58 ± 0.01 0.64 ± 0.01
(+7.8%) (-1.6%) (-28.6%) (-40.8%) (-41.8%) (-50.0%) (-48.8%)
10 UNet+ 0.56 ± 0.00 0.53 ± 0.00 0.68 ± 0.00 0.85 ± 0.00 1.00 ± 0.01 1.02 ± 0.01 1.01 ± 0.01
STCINet 0.57 ± 0.00 0.59 ± 0.00 0.74 ± 0.00 0.86 ± 0.01 1.11 ± 0.01 1.15 ± 0.01 1.26 ± 0.01
GST-UNet 0.50 ± 0.00 0.44 ± 0.01 0.45 ± 0.01 0.57 ± 0.01 0.48 ± 0.01 0.53 ± 0.01 0.49 ± 0.01
(-10.7%) (-16.7%) (-33.8%) (-32.9%) (-52.1%) (-48.0%) (-51.5%)
Figure 3: (Left) Daily PM2.5 levels across California from May to December 2018, with red lines marking major wildfires. (Center) Counties exposed to average PM2.5 $>$ 10 µg/m3 during the Camp Fire (red), origin county in dark red. (Right) Factual minus CAPO‐predicted daily respiratory admissions during peak Camp Fire. Hashed areas indicate small-population counties ($<30{,}000$).
We generate $T=200$ time steps of a $64\times 64$ ($N_{X}\times N_{Y}$) grid of observational data using the following system:
@@
$\mathbf{Y}_{t}$ $=\gamma_{0}+\gamma_{1}\bigl{(}{\color[rgb]{0.890625,0.1015625,0.109375}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.1015625,0.109375}K_{YA}*\mathbf{A}_{t-1}}\bigr{)}+\gamma_{2}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\tfrac{1}{L}\!\sum_{l=1}^{L}}\,\bigl{(}{\color[rgb]{0.0546875,0.40234375,0.890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0546875,0.40234375,0.890625}K_{YX}*\mathbf{X}_{t-l}}\bigr{)}$
$\qquad+\gamma_{3}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\mathbf{Y}_{t-1}}+\epsilon_{Y},$
where $d_{X}=1$ (one feature), “*” denotes a 2D convolution over the $N_{X}\times N_{Y}$ grid, $K_{X},K_{A},K_{YA},K_{YX}$ are $d_{k}\times d_{k}$ convolution kernels, $L$ is the number of temporal lags, and $\epsilon_{X},\epsilon_{Y}\sim\mathcal{N}(0,1)$ are i.i.d. noise terms. Each equation is evaluated at every location $s$ in the grid, so $\mathbf{X}_{t},\mathbf{A}_{t},\mathbf{Y}_{t}$ represent $N_{X}\times N_{Y}$ matrices at time $t$. We choose parameter values such that the simulation remains stable (i.e., the process does not diverge). This data-generating process (DGP) includes interference and spatial confounding from neighboring cells, as well as temporal carryover. Furthermore, $\mathbf{X}_{t}$ is a time-varying confounder, since its past values affect $\mathbf{A}$ and $\mathbf{Y}$, while current $\mathbf{A}$ influences future $\mathbf{X}$.
A concise example scenario is pollution control measures influencing health outcomes: $\mathbf{A}_{t}$ might represent binary interventions (e.g. traffic restrictions or industrial regulations), $\mathbf{X}_{t}$ could be spatially diffused air quality, and $\mathbf{Y}_{t}$ could track hospital visits or economic indicators. Government policies reacting to past pollution levels naturally create the feedback loops modeled here.
@@
5.2 Impact of Wildfires on Respiratory Health
Wildfire smoke has been associated with short-term adverse respiratory outcomes (Reid et al., 2016b, a; Cascio, 2018; Cleland et al., 2021; Letellier et al., 2025), with older adults especially vulnerable (DeFlorio-Barker et al., 2019). Recent events (e.g., ongoing Los Angeles wildfires) have amplified concerns about acute health impacts. In 2018, California experienced multiple severe wildfires (Wikipedia, 2025), including two major incidents: the Carr Fire (July–August) and the Camp Fire (November), which significantly worsened air quality across the state.
We analyze daily, county-level data from Letellier et al. (2025) that include PM2.5 (particulate matter $<\!2.5\,\mu m$), hospitalization counts for respiratory and cardiovascular conditions, and weather variables (temperature, precipitation, humidity, radiation, wind), plus population estimates from the California Department of Finance. Each of the weather variables can serve as a time-varying confounder because weather conditions affect future smoke levels and health outcomes, while also potentially being influenced by prior smoke levels.
We focus on the period from week 20 to week 48 (May 18–December 2, 2018), covering both the Carr and Camp fires. Following standard practice, we label a county as “treated” on any day it has mean PM2.5 above $10\,\mu g{/}m^{3}$, and we use the raw hospitalization counts as the outcome (rather than per-10,000 incidence to avoid instability in small counties). To represent counties in a spatial grid, we interpolate each day’s county-level data (treatment, outcome, five covariates) onto a $40\times 44$ latitude–longitude lattice (discarding cells outside California), yielding a spatiotemporal tensor of size $203\times 7\times 40\times 44$. Interpolation ensures that each grid cell approximates the region it overlaps (weighted by intersection area), so the downstream model captures spatial gradients in PM2.5, weather, and hospitalizations. We train GST-UNet with a horizon $\tau=10$. The earlier wildfire period (Jun–July) is used for validation and hyperparameter tuning. After model selection, we generate counterfactual predictions for the peak phase of the Camp Fire, November 8–17, 2018. See Appendix C for details on data preprocessing, interpolation, and masking.
@@ 6 Conclusion
We introduced GST-UNet, a neural framework for estimating causal effects in spatiotemporal settings by combining a U-Net-based architecture for spatial modeling with iterative G-computation to adjust for time-varying confounders. GST-UNet addresses key challenges such as interference, spatial confounding, temporal carryover, and time-varying confounders. Through simulations, the GST-UNet outperformed existing baselines, and in a real-world case study of wildfire smoke exposure during the 2018 Camp Fire, it provided fine-grained, location-specific, and credible effect estimates. These results highlight GST-UNet’s potential to improve causal inference in fields such as public health, environmental science, and social policy.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 1846210 and by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0023112. Part of this work was conducted while Miruna Oprescu was a research intern at Brookhaven National Laboratory.
Impact Statement
This work advances the field of machine learning by developing a spatiotemporal causal inference framework that enables more accurate estimation of treatment effects in complex real-world settings. GST-UNet has broad applications in public health, environmental science, and social policy, where understanding intervention effects can inform evidence-based decision-making. While our method is designed to improve causal inference from observational data, care must be taken when applying it to high-stakes policy decisions, ensuring robustness against biases in data collection and model assumptions. We encourage responsible use, particularly in applications affecting vulnerable populations.
@@ Appendix A Extended Literature Review
Classical Spatiotemporal Causal Inference. Early spatiotemporal causal inference methods—including spatial econometrics [Anselin, 2013], difference-in-differences [Keele and Titiunik, 2015], and synthetic controls [Ben-Michael et al., 2022]—provide useful frameworks for estimating treatment effects across regions but rely on strong assumptions such as parallel trends or stable treatment assignment. These approaches struggle with interference, nonlinear dependencies, and time-varying confounders, limiting their applicability in complex settings. More recent classical approaches estimate average treatment effects across regions or impose structural and modeling assumptions that may not generalize well to real-world spatiotemporal contexts [Wang, 2021, Christiansen et al., 2022, Papadogeorgou et al., 2022, Zhang and Ning, 2023, Zhou et al., 2024].
Machine Learning for Spatiotemporal Modeling. The rise of large-scale spatiotemporal datasets has led to the adoption of machine learning models for predictive tasks. Convolutional and recurrent neural networks [Shi et al., 2015, Zhang et al., 2017] effectively capture spatial and temporal dependencies but lack explicit causal adjustment mechanisms. Graph-based deep learning methods [Li et al., 2017, Wu et al., 2019] model spatial interactions but typically ignore feedback effects from time-varying confounders. Some recent work integrates spatial representations for causal inference—e.g., Tec et al. [2023] incorporate geographic confounders using a UNet-based model—but these methods do not explicitly model iterative dependencies over time or adjust for time-varying confounders.
Time-Series Causal Inference. In the longitudinal domain, time-series causal inference typically employs recurrent networks, Transformers, or propensity-based models [Bica et al., 2020, Seedat et al., 2022, Melnychuk et al., 2022], but these approaches often assume independent time series, overlooking spatial interference and cross-series confounding. Handling time-varying confounders has relied on marginal structural models [Robins et al., 2000] or iterative G-computation [Bang and Robins, 2005, Robins and Hernan, 2008], but machine learning adaptations [Lim, 2018, Li et al., 2021, Hess et al., 2024] continue to assume independent observations, making them ill-suited for fully spatiotemporal settings where interference is prevalent.
Neural-Based Spatiotemporal Causal Inference. Recent efforts have explored neural models for spatiotemporal causal inference. Tec et al. [2023] integrate spatial confounding adjustments using a UNet-based framework but do not model feedback effects from time-varying confounders. The most similar work to ours, Ali et al. [2024], introduces a climate-focused neural model that shares certain architectural components but is designed primarily for predictive tasks rather than causal identification, leaving time-varying confounders unaddressed.
Our Contribution. GST-UNet bridges these gaps by combining a U-Net-based architecture for spatiotemporal grids with iterative G-computation to adjust for time-varying confounders. Unlike prior methods, GST-UNet explicitly accounts for interference, nonlinear spatial-temporal dynamics, and feedback loops, ensuring valid causal identification under standard assumptions. This enables fine-grained, location-specific estimates of potential outcomes, improving spatiotemporal causal inference in domains where observational data is abundant but controlled experiments are infeasible.
Appendix B Proof of Theorem 3.3

### v2

@@ 6 Experiments
Additional details–including exact simulation parameters, model architecture and execution setups, hyperparameter selection strategies, and validation procedures–can be found in Appendix˜D. Replication code is available at https://github.com/moprescu/GSTUNet.
@@ 6.1 Synthetic Data
We generate $T=200$ time steps of a $64\times 64$ ($N_{X}\times N_{Y}$) grid of observational data using the following data generating process (DGP):
@@
$\mathbf{Y}_{t}$ $=\gamma_{0}+\gamma_{1}\bigl({\color[rgb]{0.890625,0.1015625,0.109375}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.1015625,0.109375}K_{YA}*\mathbf{A}_{t-1}}\bigr)+\gamma_{2}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\tfrac{1}{L}\!\sum_{l=1}^{L}}\,\bigl({\color[rgb]{0.0546875,0.40234375,0.890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.0546875,0.40234375,0.890625}K_{YX}*\mathbf{X}_{t-l}}\bigr)+\gamma_{3}{\color[rgb]{0.13671875,0.546875,0.26953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.546875,0.26953125}\mathbf{Y}_{t-1}}+\epsilon_{Y},$
where $d_{X}=1$, "*" denotes a $3{\times}3$ spatial convolution over the $N_{X}\times N_{Y}$ grid, and $\epsilon_{X},\epsilon_{Y}\sim\mathcal{N}(0,1)$ are i.i.d. noise. Each kernel $K_{X},K_{A},K_{YA},K_{YX}$ encodes a local advection–diffusion process that mimics wind-driven pollutant transport, with interventions $\mathbf{A}_{t}$ injecting additional emissions that propagate through the same kernel. This physically realistic setup produces interference, spatial confounding, and temporal carryover—the three challenges GST-UNet is designed to address. Each equation is evaluated at every spatial location, so $\mathbf{X}_{t}$, $\mathbf{A}_{t}$, and $\mathbf{Y}_{t}$ are $N_{X}\times N_{Y}$ matrices. Here, $\mathbf{X}_{t}$ acts as a time-varying confounder: its past influences both $\mathbf{A}_{t}$ and $\mathbf{Y}_{t}$, while current interventions $\mathbf{A}_{t}$ affect future $\mathbf{X}_{t+1}$. For example, $\mathbf{A}_{t}$ may represent regulatory actions, $\mathbf{X}_{t}$ air quality, and $\mathbf{Y}_{t}$ health outcomes—capturing feedback from policy to exposure to outcome, and back to future policy.
Table 1: RMSE $\pm$ SD across test trajectories. Bold indicates lowest error per column; color shows improvement (RMSE decrease or increase) over best baseline (excluding ablations).
$\tau$ Model $\beta_{1}=0.0$ $\beta_{1}=0.5$ $\beta_{1}=1.0$ $\beta_{1}=1.5$ $\beta_{1}=2.0$
5 UNet+ 0.28 ± 0.00 0.36 ± 0.00 0.54 ± 0.01 0.71 ± 0.01 0.81 ± 0.01
STCINet 0.29 ± 0.00 0.38 ± 0.01 0.62 ± 0.01 0.80 ± 0.01 0.90 ± 0.01
IPWUNet 0.60 ± 0.01 0.58 ± 0.01 0.58 ± 0.01 0.59 ± 0.01 0.59 ± 0.01
GST-UNet w/o Attention 0.50 ± 0.00 0.46 ± 0.00 0.51 ± 0.00 0.45 ± 0.01 0.47 ± 0.01
GST-UNet w/o Curriculum 0.69 ± 0.00 0.64 ± 0.00 0.63 ± 0.00 0.61 ± 0.01 0.61 ± 0.01
GST-UNet 0.33 ± 0.00 0.35 ± 0.00 0.40 ± 0.00 0.44 ± 0.00 0.40 ± 0.01
(+17.9%) (-2.7%) (-21.6%) (-25.4%) (-32.2%)
10 UNet+ 0.28 ± 0.00 0.61 ± 0.00 1.18 ± 0.00 1.45 ± 0.00 1.71 ± 0.01
STCINet 0.31 ± 0.00 0.68 ± 0.00 1.25 ± 0.00 1.47 ± 0.01 1.60 ± 0.01
IPWUNet 0.78 ± 0.01 0.80 ± 0.01 0.96 ± 0.01 1.19 ± 0.02 1.08 ± 0.01
GST-UNet w/o Attention 0.42 ± 0.00 0.60 ± 0.00 0.61 ± 0.00 0.79 ± 0.01 1.07 ± 0.01
GST-UNet w/o Curriculum 0.62 ± 0.00 0.88 ± 0.00 1.02 ± 0.00 1.08 ± 0.01 1.12 ± 0.01
GST-UNet 0.38 ± 0.00 0.55 ± 0.00 0.68 ± 0.00 0.73 ± 0.01 0.85 ± 0.01
(+35.7%) (-9.8%) (-29.2%) (-38.7%) (-21.3%)
We vary $\beta_{1}$ to control time-varying confounding: when $\beta_{1}=0$, $\mathbf{X}_{t}$ does not affect $\mathbf{A}_{t}$, eliminating confounding; larger values increase its strength. For each $\beta_{1}$, we generate 50 test trajectories from random initial states, fix their histories, and simulate 100 $\tau$-step counterfactual futures to estimate true CAPOs, with $\tau\in\{5,10\}$. We compare GST-UNet against three baselines: (i) UNet+, which uses a U-Net + ConvLSTM + Attention backbone with $A_{t}$ as an input channel but performs no iterative adjustment; (ii) STCINet [1], which estimates direct and indirect effects without modeling time-varying confounding; and (iii) IPWUNet, an inverse-propensity-weighting variant that reweights pseudo-outcomes using a UNet-style propensity estimator but cannot correct for spatial interference (details in Appendix˜D). We also test ablations of GST-UNet without curriculum or attention. Table˜1 shows that when $\beta_{1}=0$, UNet+ performs best—G-computation is unnecessary and adds noise. As $\beta_{1}$ increases, UNet+ and STCINet degrade sharply, while GST-UNet remains stable. IPWUNet shows some benefit but is biased even at $\beta_{1}=0$ due to uncorrected interference. GST-UNet consistently outperforms all baselines, demonstrating the value of iterative G-computation. Curriculum training substantially improves performance across horizons, while attention yields modest gains—consistent with our predominantly local dynamics. Additional ablation analyses, including neighbor aggregation experiments, are reported in Appendix˜D.
6.2 Impact of Wildfires on Respiratory Health
Wildfire smoke has been linked to short-term respiratory harms [34, 33, 10, 12, 23], with older adults especially vulnerable [14]. At the time this work was conducted (January 2025), a series of 14 destructive wildfires affected the Los Angeles metropolitan area and San Diego County in California, underscoring the urgency of understanding the health impacts of such events. In this study, we focus on a previous large-scale episode: the 2018 California wildfire season [45], which included the Carr Fire (July–August) and the Camp Fire (November) and significantly worsened air quality.
We use daily, county-level data from Letellier et al. [23] (see Section˜D.2), including PM2.5, respiratory/cardiovascular hospitalizations, and weather variables (temperature, precipitation, humidity, radiation, wind), along with population estimates from the California Department of Finance. Each of the weather variables can be a time-varying confounder: weather conditions affect future smoke levels and health outcomes, while also being influenced by prior smoke levels.
@@ 6.2 Impact of Wildfires on Respiratory Health
Figure 3: (Left) Daily PM2.5 levels across California from May to December 2018, with red lines marking major wildfires. (Center) Counties exposed to average PM2.5 >> 10 µg/m3 during the Camp Fire (red), origin county in dark red. (Right) Factual minus CAPO‐predicted daily respiratory admissions during peak Camp Fire. Hashed areas indicate small-population counties (<30,000<30{,}000).
@@ 7 Conclusion
We presented GST-UNet, a neural framework for spatiotemporal causal inference that combines U-Net–based representation learning with iterative G-computation to adjust for time-varying confounders. GST-UNet addresses key challenges such as interference, spatial confounding, temporal carryover, and time-varying feedback. We establish theoretical identification and consistency guarantees, validate performance in synthetic settings with controlled confounding, and demonstrate practical utility in estimating the impact of wildfire smoke exposure during the 2018 Camp Fire. Together, these results position GST-UNet as a ready-to-use tool for practitioners, offering reliable, interpretable causal estimates in complex spatiotemporal environments. We discuss limitations and broader impacts in Appendix˜E.
@@ Acknowledgements
We thank the anonymous reviewers for their thoughtful feedback and the constructive dialogue during the review process, which greatly clarified our exposition and strengthened the final version of this work. Miruna Oprescu and BNL team (D. Park, X. Luo, S. Yoo) were supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Awards DE-SC0023112 and DE-SC0012704, respectively. Nathan Kallus was supported by the U.S. National Science Foundation under Grant No. 1846210. Part of this work was conducted while Miruna Oprescu was a research intern at Brookhaven National Laboratory. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. Department of Energy or the U.S. National Science Foundation.
@@ Appendix A Extended Literature Review
Table 2: Key differences between prior neural G-computation methods and GST-UNet.
Aspect Prior Neural G-Computation GST-UNet (ours)
Data structure Many independent temporal trajectories (e.g., patient sequences); no inter-unit interactions such as spatial dependence. Single spatiotemporal chain where outcomes, covariates, and treatments evolve jointly across a lattice; strong spatial coupling and interference.
Encoder RNN/Transformer over time only. ConvLSTM-UNet encoder aggregates neighbour covariates/treatments before G-heads, capturing interference and spatial confounding.
Training Standard end-to-end due to i.i.d trajectories; stability arises from large data rather than curriculum or spatial priors. Curriculum-stabilized multi-head training for accurate pseudo-outcome generation under limited samples.
Theory Classical G-formula under i.i.d. trajectories; no single-chain guarantees. Identification (Theorem˜1) and consistency (Theorem˜3) under representation-based time-invariance for a single chain.
@@
Classical Spatiotemporal Causal Inference. Early spatiotemporal causal inference methods–including spatial econometrics [2], difference-in-differences [20], and synthetic controls [4]–provide useful frameworks for estimating treatment effects across regions but rely on strong assumptions such as parallel trends or stable treatment assignment. These approaches struggle with interference, nonlinear dependencies, and time-varying confounders, limiting their applicability in complex settings. More recent approaches for spatiotemporal causal inference handle time-varying confounding through inverse propensity weighting (IPW), typically by extending marginal structural models to the spatial or spatiotemporal domain. For instance, Papadogeorgou et al. [31] and Zhou et al. [49] employ IPW-style adjustments to estimate regional average treatment effects across space and time. However, these approaches cannot accommodate interference unless strong assumptions are made–e.g., defining a user-specified exposure mapping or restricting attention to hyper–local interactions (see also [44, 11, 31, 48]). Such simplifications may be ill-suited for real-world systems with rich spatial dependencies. Moreover, even recent advances in this space remain limited; as noted by Zhou et al. [49], the literature on spatiotemporal causal inference remains sparse, especially in settings with feedback loops or time-varying confounding.
Machine Learning for Spatiotemporal Modeling. Spatiotemporal predictive modeling has seen rapid progress with the rise of deep learning. Convolutional and recurrent neural networks are widely used for forecasting spatially indexed time series (e.g., weather or traffic) [40, 47], while graph-based methods (e.g., Graph WaveNet [46], Diffusion Convolutional RNN [25]) capture non-Euclidean spatial dependencies. Vision transformer variants, including Video Swin Transformers [27] and TimeSformer [6], extend attention-based models to spatiotemporal video data. These architectures can learn complex non-local interactions over space and time. However, such models are typically optimized for prediction tasks and do not include causal adjustments. Without mechanisms like propensity modeling or G-computation, they remain ill-equipped to estimate counterfactual outcomes or adjust for time-varying confounding. Some recent work integrates spatial representations for causal inference–e.g., Tec et al. [42] incorporate non-local confounders using a UNet-based model–but these methods do not explicitly model dependencies over time or adjust for time-varying confounders.
Time-Series Causal Inference. In the longitudinal domain, time-series causal inference has developed tools for handling temporal confounding using models such as marginal structural models [36], IPW-style estimation [26], and iterative G-computation [35]. Recent ML-based extensions include recurrent networks [7, 24, 39], Transformers [28, 18] and meta-learners [16]. However, all these methods assume access to independent time series–e.g., across units or patients–which allows for pooling across trajectories. These methods do not consider spatial dependencies, interference, or scenarios with a single observed spatiotemporal realization. As such, while they may handle time-varying confounding, they do not generalize to our setting. Table˜2 summarizes the key methodological differences between GST-UNet and prior neural G-computation frameworks.
Neural-Based Spatiotemporal Causal Inference. There has been limited work on neural models that explicitly address spatiotemporal causal inference. Tec et al. [42] use a U-Net backbone to learn spatial representations for causal inference in air pollution studies but do not address time-varying confounding or feedback loops. Ali et al. [1] present a U-Net–based architecture for predicting direct and indirect effects in climate contexts, but primarily focus on forecasting rather than causal identification. While these works highlight growing interest in neural approaches to causal inference in spatiotemporal domains, none incorporate an iterative adjustment procedure like G-computation that handles time-varying confounders, leaving identification in these settings largely unaddressed.
Our Contribution. GST-UNet bridges these gaps by combining flexible spatiotemporal neural architectures with a theoretically grounded iterative G-computation framework. This allows valid estimation of potential outcomes in the presence of interference, spatial confounding, and time-varying confounding–without requiring practitioners to specify structural models or exposure mappings. To our knowledge, this is the first end-to-end framework to implement G-computation for causal inference over a single spatiotemporal trajectory. We integrate spatiotemporal processing via U-Nets and ConvLSTMs with a principled multi-head neural causal module, and we design a curriculum-based training strategy to stabilize learning of recursive pseudo-outcomes. Together, these components yield a ready-to-use tool for practitioners, with consistent identification guarantees and robust empirical performance. By abstracting away the modeling choices typically required in structural spatiotemporal methods, GST-UNet makes spatiotemporal causal estimation more accessible, interpretable, and reliable for real-world applications.
Appendix B Proof of Theorem 1